OpenStudy (anonymous):

z - (y − y); use y = 6, and z = 2

OpenStudy (anonymous):

z - (y − y) =z - 0 =z=2

OpenStudy (anonymous):

Since y-y =0

OpenStudy (anonymous):

Ok thanks, that helps a lot